# Proton mass decompositions in the NNLO QCD

### v1

@@ Abstract
The invariant mass of (2) leads to the well-known proton mass formula [21] associated with the trace anomaly [6, 7, 8], i.e., $T^{\alpha}_{\ \alpha}=\frac{\beta(g)}{2g}F^{2}+[1+\gamma_{m}(g)]m\bar{\psi}\psi$, with the QCD $\beta$-function $\beta(g)$ and the anomalous dimension for the quark mass, $\gamma_{m}(g)$. Thus, the proton mass is expressed by the non-perturbative matrix elements $\langle p|F^{2}|p\rangle$, $\langle p|m\bar{\psi}\psi|p\rangle$; these matrix elements of the Lorentz scalar operators equal the corresponding matrix elements at the rest frame, as $\langle p|{\cal O}(0)|p\rangle$ $\langle p|{\cal O}(0)|p\rangle|_{\bm{p}=\bm{0}}=2M\langle\int d^{3}x{\cal O}(x)\rangle$, with ${\cal O}(x)=F^{2}(x),m\bar{\psi}\psi(x)$, where we have introduced the shorthand notation for the expectation value in the rest frame as $\langle\,\cdots\,\rangle\equiv(\langle p|\cdots|p\rangle/\langle p|p\rangle)_{\bm{p}=\bm{0}}$. As discussed below, $\langle p|m\bar{\psi}\psi|p\rangle$ is expressed by the sigma terms determined (semi)emprically; then, the condition that the second equation of (2) should reproduce the value of the proton mass allows us to evaluate the proton matrix element $\langle p|F^{2}|p\rangle$. Now, using those non-perturbative values of $\langle p|F^{2}|p\rangle$ as well as of $\langle p|m\bar{\psi}\psi|p\rangle$, the separation of $T^{\alpha\beta}$ into (1) allows us to decompose the proton mass into the quark and gluon contributons as $M={\cal M}_{q}+{\cal M}_{g}$ (see (14) below), combined with the trace anomaly, as taking trace after renormalization  [^2] Taking trace and performing renormalization do not commute; if their order were reversed, as taking trace before renormalization, the RHS of (3) would be modified into $m\bar{\psi}\psi$, which corresponds to the “D2 scheme”[14, 17, 15]., for each part of (1),
@@
and $T_{g\ \alpha}^{\alpha}=T^{\alpha}_{\ \alpha}-T_{q\ \alpha}^{\alpha}$, with $\alpha_{s}\equiv g^{2}/4\pi$, for $N_{c}$ color with $C_{F}=(N_{c}^{2}-1)/2N_{c}$; (3) holds in the MS-like schemes, and the ellipses stand for the corrections at the two loops (${\cal O}(\alpha_{s}^{2})$) and higher, which are derived in [5, 9, 10]. The quantitative results using the three-loop formulas [9] for (3) will be discussed in Figs. 2, 4 below.
@@
where $\overline{p}^{\alpha}\equiv(p^{\alpha}+p^{\prime\alpha})/2$, $u(p)$ is the proton spinor, and the ellipses stand for the terms associated with the other Lorentz structures that vanish as $\Delta=p^{\prime}-p\to 0$ (see e.g., [18, 23, 24, 25]). The sum rule, $\bar{C}_{q}(\Delta^{2},\mu)+\bar{C}_{g}(\Delta^{2},\mu)=0$, is derived by contracting (5) with $\Delta_{\alpha}$ and using $\partial_{\alpha}T^{\alpha\beta}=0$. Also, $\langle p|T^{\alpha\beta}|p\rangle=2p^{\alpha}p^{\beta}$ ensures the forward sum rule, $A_{q}(0,\mu)+A_{g}(0,\mu)=1$. Physically, the gravitational form factors represent the coupling of the proton with a graviton. Instead of probing them via such impractical graviton-proton processes, the gravitational form factors can be accessed through the spin-2 contributions induced by the two vector currents; e.g., with the two electromagnetic currents, the deeply virtual Compton scatterings [26] in JLab [27] and the future electron-ion collider (EIC) allow us to probe the gravitational form factors. Substituting the forward limit of (5) into $M=\langle\,H\,\rangle$, we obtain the “two-term decomposition” [13] using the $\Delta=0$ values of the gravitational form factors as
@@
The values of $\bar{C}_{r}(0,\mu)$ are also available with the same level accuracy as in $A_{r}(0,\mu)$: the trace of both sides of (5) allows us to write down the formula [5],
@@
with $M_{q}=U_{q}-M_{m}-({\cal M}_{q}-M_{m})/4$, $M_{g}=U_{g}-{\cal M}_{g}/4$; this corresponds to the famous “four-term decomposition” proposed by Ji [11, 4, 12] and coincides with Eq.(62) of [17]. For the $n_{f}=3$ active quark flavors, $M_{m}$ of (10) can be expressed as $M_{m}=\langle p|[(m_{u}+m_{d})/2](\bar{u}u+\bar{d}d)+m_{s}\bar{s}s|p\rangle/2M=\sigma_{\pi N}+\sigma_{s}$, up to small isospin-violating corrections of ${\cal O}(m_{d}-m_{u})$, using the pion-nucleon sigma-term $\sigma_{\pi N}$ and the strangeness content $\sigma_{s}$. Here, we use $\sigma_{\pi N}=0.0591(35)$ GeV due to a recent phenomenological analysis [29] and $\sigma_{s}=0.0456(62)$ GeV due to a recent lattice QCD determination [30]; these non-perturbative matrix elements arise also in (8) with (3), and the above-mentioned NNLO values of $\bar{C}_{q,g}(0,\mu)$ (e.g., (9)) are calculated using the same input values for $\sigma_{\pi N},\sigma_{s}$.
@@
The composite operators are classified by the twist that equals their “dimension minus spin”, and the traceless and trace parts of (1) are of twist two and four, respectively. Comparing this with (5), $A_{r}(0,\mu)$ is of twist two, and this implies (iii) above. The studies of hard processes like the deep inelastic lepton-proton scatterings (DIS) clarified [19] that the twist-two operators in QCD are associated with the independent partonic motion inside the proton, according to the asymptotic freedom, while the non-perturbative QCD interactions induce the correlations among the partons, emerging as matrix elements of the operators of twist four and higher  [^6] For the quantities dependent on the polarization, the twist-three operators also play roles to describe the spin-dependent partonic correlations, see e.g., [19, 34, 35, 36].. As noted in (iii), $M_{q}$ does not match with such physical meaning of the twist, although $M_{q}$ is intended [11, 4] to represent the contribution of the quark’s gauge-invariant “kinetic and potential energies” operator $\bar{\psi}i\bm{\gamma}\cdot\overleftrightarrow{\bm{D}}\psi$ of (4). The physical meaning of the twist calls for the rearrangement of the conventional decomposition (12), such that the terms of different twist in each of the quark and gluon contributions are completely separated (see the formulas in (iii)): $M_{q}+M_{m}=(M_{q}+3M_{m}/4)+M_{m}/4=3MA_{q}(0,\mu)/4+M_{m}/4$, where $M_{m}/4$ is to be combined with (11), leads to the new four-term decomposition,
@@
As explained in [20], the global QCD analyses for the pion’s PDFs are available at the NLO level, and the resultant value, $A_{q}^{\pi}(0,\mu_{0})=0.58(9)$ [40], has rather large ($\sim$ ten percent level) uncertainties (see also [41, 42, 43]), while chiral perturbation theory gives the accurate value for $M_{m}^{\pi}=\langle\pi|m\bar{\psi}\psi|\pi\rangle/2M^{\pi}$ corresponding to (10): $M_{m}^{\pi}=M^{\pi}/2$, to the $O(6\%)$ correction [44, 45]. Using these input values, $A_{q,g}^{\pi}(0,\mu)$, $\bar{C}^{\pi}_{q,g}(0,\mu)$ have been calculated [20] by our NNLO formulas in the $\overline{\rm MS}$ scheme as, e.g., $\bar{C}^{\pi}_{q}(0,\mu_{0})|_{{n_{f}}=3}=-\bar{C}^{\pi}_{g}(0,\mu_{0})|_{{n_{f}}=3}=-0.03(2)$ with $\mu_{0}=2$ GeV. These values allow us to evaluate the pion-mass decompositions corresponding to the four-term formulas (12) and (13), as displayed in Figs. 3 and 4, respectively, where the presented results have the uncertainties of $\sim$ ten percent level; we recognize that, also in the pion case, the points (i)-(iii) discussed above are improved using (13). The comparison between Figs. 2 and 4 reveals: the twist-two contributions behave similarly, while the proton and the pion are discriminated by the behaviors of the twist-four components, as driven by the large quark-mass term $M_{m}^{\pi}$ for the pion, which reflects the Nambu-Goldstone boson nature. Such difference could not be observed, if the last two terms of twist four in (13) were not subdivided into the quark and gluon contributions, because ${\cal M}_{q}/4+{\cal M}_{g}/4=M/4$; on the other hand, such difference could be observed, when the “D2 scheme” [14, 17] is employed as in [15] for the twist-four contributions in (13) by making the replacements ${\cal M}_{q}\to M_{m}$, ${\cal M}_{g}\to M_{a}$, but we stick to the $\overline{\rm MS}$ scheme to treat the twist-four effects as well as twist-two effects for our quantitative evaluation.

### v2

@@ Abstract
The invariant mass of (2) leads to the well-known proton mass formula [21] associated with the trace anomaly [6, 7, 8], i.e., $T^{\alpha}_{\ \alpha}=\frac{\beta(g)}{2g}F^{2}+[1+\gamma_{m}(g)]m\bar{\psi}\psi$, with the QCD $\beta$-function $\beta(g)$ and the anomalous dimension for the quark mass, $\gamma_{m}(g)$. Thus, the proton mass is expressed by the non-perturbative matrix elements $\langle p|F^{2}|p\rangle$, $\langle p|m\bar{\psi}\psi|p\rangle$; these matrix elements of the Lorentz scalar operators equal the corresponding matrix elements at the rest frame, as $\langle p|O(0)|p\rangle$ $\langle p|O(0)|p\rangle|_{\bm{p}=\bm{0}}=2M\langle\int d^{3}xO(x)\rangle$, with $O(x)=F^{2}(x),m\bar{\psi}\psi(x)$, where we have introduced the shorthand notation for the expectation value in the rest frame as $\langle\,\cdots\,\rangle\equiv(\langle p|\cdots|p\rangle/\langle p|p\rangle)_{\bm{p}=\bm{0}}$. As discussed below, $\langle p|m\bar{\psi}\psi|p\rangle$ is expressed by the sigma terms determined (semi)emprically; then, the condition that the second equation of (2) should reproduce the value of the proton mass allows us to evaluate the proton matrix element $\langle p|F^{2}|p\rangle$. Now, using those non-perturbative values of $\langle p|F^{2}|p\rangle$ as well as of $\langle p|m\bar{\psi}\psi|p\rangle$, the separation of $T^{\alpha\beta}$ into (1) allows us to decompose the proton mass into the quark and gluon contributons as $M={\cal M}_{q}+{\cal M}_{g}$ (see (14) below), combined with the trace anomaly, as taking trace after renormalization  [^2] Taking trace and performing renormalization do not commute; if their order were reversed, as taking trace before renormalization, the RHS of (3) would be modified into $m\bar{\psi}\psi$, which corresponds to the “D2 scheme”[14, 17, 15]., for each part of (1),
@@
and $T_{g\ \alpha}^{\alpha}=T^{\alpha}_{\ \alpha}-T_{q\ \alpha}^{\alpha}$, with $\alpha_{s}\equiv g^{2}/4\pi$, for $n_{f}$ flavor and $N_{c}$ color with $C_{F}=(N_{c}^{2}-1)/2N_{c}$; (3) holds in the MS-like schemes, and the ellipses stand for the corrections at the two loops (${\cal O}(\alpha_{s}^{2})$) and higher, which are derived in [5, 9, 10]. The quantitative results using the three-loop formulas [9] for (3) will be discussed in Figs. 2, 4 below.
@@
where $\overline{p}^{\alpha}\equiv(p^{\alpha}+p^{\prime\alpha})/2$, $u(p)$ is the proton spinor, and the ellipses stand for the terms associated with the other Lorentz structures that vanish as $\Delta=p^{\prime}-p\to 0$ (see e.g., [18, 23, 24, 25]). The sum rule, $\bar{C}_{q}(\Delta^{2},\mu)+\bar{C}_{g}(\Delta^{2},\mu)=0$, is derived by contracting (5) with $\Delta_{\alpha}$ and using $\partial_{\alpha}T^{\alpha\beta}=0$. Also, $\langle p|T^{\alpha\beta}|p\rangle=2p^{\alpha}p^{\beta}$ ensures the forward sum rule, $A_{q}(0,\mu)+A_{g}(0,\mu)=1$. Physically, the gravitational form factors represent the coupling of the proton with a graviton. Instead of probing them via such impractical graviton-proton processes, the gravitational form factors can be accessed through the spin-2 contributions induced by the two vector currents; e.g., with the two electromagnetic currents, the deeply virtual Compton scatterings [26] in JLab [27] and the future electron-ion collider (EIC) [28] allow us to probe the gravitational form factors. Substituting the forward limit of (5) into $M=\langle\,H\,\rangle$, we obtain the “two-term decomposition” [13] using the $\Delta=0$ values of the gravitational form factors as
@@
The values of $\bar{C}_{r}(0,\mu)$ are also available with the same level of accuracy as in $A_{r}(0,\mu)$: the trace of both sides of (5) allows us to write down the formula [5],
@@
with $M_{q}=U_{q}-M_{m}-({\cal M}_{q}-M_{m})/4$, $M_{g}=U_{g}-{\cal M}_{g}/4$; this corresponds to the famous “four-term decomposition” proposed by Ji [11, 4, 12] and coincides with Eq.(62) of [17]. For the $n_{f}=3$ active quark flavors, $M_{m}$ of (10) can be expressed as $M_{m}=\langle p|[(m_{u}+m_{d})/2](\bar{u}u+\bar{d}d)+m_{s}\bar{s}s|p\rangle/2M=\sigma_{\pi N}+\sigma_{s}$, up to small isospin-violating corrections of ${\cal O}(m_{d}-m_{u})$, using the pion-nucleon sigma-term $\sigma_{\pi N}$ and the strangeness content $\sigma_{s}$. Here, we use $\sigma_{\pi N}=0.0591(35)$ GeV due to a recent phenomenological analysis [30] and $\sigma_{s}=0.0456(62)$ GeV due to a recent lattice QCD determination [31]; see also [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. These non-perturbative matrix elements arise also in (8) with (3), and the above-mentioned NNLO values of $\bar{C}_{q,g}(0,\mu)$ (e.g., (9)) are calculated using the same input values for $\sigma_{\pi N},\sigma_{s}$.
@@
The composite operators are classified by the twist that equals their “dimension minus spin”, and the traceless and trace parts of (1) are of twist two and four, respectively. Comparing this with (5), $A_{r}(0,\mu)$ is of twist two, and this implies (iii) above. The studies of hard processes like the deep inelastic lepton-proton scatterings clarified [19] that the twist-two operators in QCD are associated with the independent partonic motion inside the proton, according to the asymptotic freedom, while the non-perturbative QCD interactions induce the correlations among the partons, emerging as matrix elements of the operators of twist four and higher  [^6] For the quantities dependent on the polarization, the twist-three operators also play roles to describe the spin-dependent partonic correlations, see e.g., [19, 46, 47, 48].. As noted in (iii), $M_{q}$ does not match with such physical meaning of the twist, although $M_{q}$ is intended [11, 4] to represent the contribution of the quark’s gauge-invariant “kinetic and potential energies” operator $\bar{\psi}i\bm{\gamma}\cdot\overleftrightarrow{\bm{D}}\psi$ of (4). The physical meaning of the twist calls for the rearrangement of the conventional decomposition (12), such that the terms of different twist in each of the quark and gluon contributions are completely separated (see the formulas in (iii)): $M_{q}+M_{m}=(M_{q}+3M_{m}/4)+M_{m}/4=3MA_{q}(0,\mu)/4+M_{m}/4$, where $M_{m}/4$ is to be combined with (11), leads to the new four-term decomposition,
@@
As explained in [20], the global QCD analyses for the pion’s PDFs are available at the NLO level, and the resultant value, $A_{q}^{\pi}(0,\mu_{0})=0.58(9)$ [52], has rather large ($\sim$ ten percent level) uncertainties (see also [53, 54, 55]), while chiral perturbation theory gives the accurate value for $M_{m}^{\pi}=\langle\pi|m\bar{\psi}\psi|\pi\rangle/2M^{\pi}$ corresponding to (10): $M_{m}^{\pi}=M^{\pi}/2$, to the ${\cal O}(6\%)$ correction [56, 57]. Using these input values, $A_{r}^{\pi}(0,\mu)$, $\bar{C}^{\pi}_{r}(0,\mu)$ have been calculated [20] by our NNLO formulas in the $\overline{\rm MS}$ scheme as, e.g., $\bar{C}^{\pi}_{q}(0,\mu_{0})|_{{n_{f}}=3}=-\bar{C}^{\pi}_{g}(0,\mu_{0})|_{{n_{f}}=3}=-0.03(2)$ with $\mu_{0}=2$ GeV. These values allow us to evaluate the pion-mass decompositions corresponding to the four-term formulas (12) and (13), as displayed in Figs. 3 and 4, respectively, where the presented results have the uncertainties of $\sim$ ten percent level; we recognize that, also in the pion case, the points (i)-(iii) discussed above are improved using (13). The comparison between Figs. 2 and 4 reveals: the twist-two contributions behave similarly, while the proton and the pion are discriminated by the behaviors of the twist-four components, as driven by the large quark-mass term $M_{m}^{\pi}$ for the pion, which reflects the Nambu-Goldstone boson nature. Such difference could not be observed, if the last two terms of twist four in (13) were not subdivided into the quark and gluon contributions, because ${\cal M}_{q}/4+{\cal M}_{g}/4=M/4$; on the other hand, such difference could be observed, when the “D2 scheme” [14, 17] is employed as in [15] for the twist-four contributions in (13) by making the replacements ${\cal M}_{q}\to M_{m}$, ${\cal M}_{g}\to M_{a}$, but we stick to the $\overline{\rm MS}$ scheme to treat the twist-four effects as well as twist-two effects for our quantitative evaluation.